The angle of depression from the top of a building to apoint on the groudn is 32 deg 30'. How far is the point on the ground from the top of the building if the buildlng is 252 m high?
ashley917@netzero.net

Answer by stanbon(75887)

The angle of depression from the top of a building to apoint on the groudn is 32 deg 30'. How far is the point on the ground from the top of the building if the buildlng is 252 m high?
Draw the picture.
You have a right triangle with an acute angle of 32 deg 30' and
a side opposite the acute angle of length 252 m.
Use the sine to find the hypotenuse, as follows:
sin(32 deg 30')=252/hyp.
hyp=252/sin(32 deg 30')
hyp=252/0.53729961...
hyp=469 m
Cheers,
Stan H.